OpenStudy (anonymous):

The 8.00-cm long second hand on a watch rotates smoothly. What is its angular velocity? What is the speed of the tip of the hand?

OpenStudy (mayankdevnani):

@iphan7 is angular velocity be given ?

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

what formular do you use? r= 8.0 cm is it right?

OpenStudy (mayankdevnani):

okk, lets assume that angular velocity be ω

OpenStudy (mayankdevnani):

and radius is 8 cm

OpenStudy (mayankdevnani):

 now use :- \[\large \bf v=\omega \times r\]

OpenStudy (mayankdevnani):

understood? @Iphan7

OpenStudy (anonymous):

so there is only 1 number is given how can I calculate?

OpenStudy (mayankdevnani):

\[\large \bf v= \omega \times r\] \[\large \bf v= \omega \times 8=8 \omega=\color{red}{Answer}\]

OpenStudy (anonymous):

the answer need to be in rad/s

OpenStudy (mayankdevnani):

see, second hand moves \(\large \bf 2 \pi r\) to complete one revolution in 60s. therefore, \[\large \bf \omega=\frac{2 \pi r}{60}=\frac{2 \pi \times 0.08}{60}=0.0026 \pi ~rad/s\]

OpenStudy (anonymous):

I got it. I have to calculate for the V by calculate the perimeter of circle 1st

OpenStudy (mayankdevnani):

yes !! (mentioned above)

OpenStudy (anonymous):

thanks for reponsed

OpenStudy (mayankdevnani):

!!!!